Theory and Applications of Categories, Vol. 8, No. 4, 2001, pp. 54–62.

ON MACKEY TOPOLOGIES IN TOPOLOGICAL ABELIAN GROUPS

MICHAEL BARR AND HEINRICH KLEISLI

ABSTRACT. Let C be a full subcategory of the category of topological abelian groups and SPC denote the full subcategory of subobjects of products of objects of C. We say that SPC has Mackey coreflections if there is a functor that assigns to each object A of SPC an object τA that has the same group of characters as A and is the finest topology with that property. We show that the existence of Mackey coreflections in SPC is equivalent to the injectivity of the circle with respect to topological subgroups of groups in C.

1. Introduction

Consider a full subcategory C of the category of topological abelian groups that includes the circle group T = R/Z. If A is a topological group, then a character on A is a continuous homomorphism χ: A ⟶ T. We let A denote the discrete group of all characters of A. Let SPC denote the closure of C with respect to arbitrary products and subobjects.

If A is an object of SPC, we say that A is a Mackey group in SPC if, whenever Ã in SPC is the same underlying group as A with a topology finer than that of A such that the identity Ã ⟶ A induces an isomorphism A ⟶ Ã , then Ã = A. In other words, if A has the finest possible topology for an object of SPC with the same elements and the same group of characters. We denote by SPC_{τ} the full subcategory of Mackey groups.

We say that SPC admits Mackey coreflections if there is a functor τ: SPC ⟶ SPC that assigns to each group A of SPC a Mackey group τA that has the same underlying group group as A as well as the same set of functionals and additionally has the property that for f: A ⟶ B, then τf: τA ⟶ τB is the same function as f.

Since τA has the finest topology among the objects of SPC that have the same elements and the same characters as A it follows that the topology of τA is finer than that of A, so that the identity function τA ⟶ A is continuous. Since for f: A ⟶ B, τf = f, it follows that

[figure: A; B; f; τA; τB; τf; =; =]

The first author would like to acknowledge the support of the NSERC of Canada and both authors would like to acknowledge support of the Swiss National Science Foundation (Project 2000-056811.99).

Received by the editors 2000 July 13 and, in revised form, 2000 September 9.

Transmitted by Walter Tholen. Published on 2001 February 5. See note on p.62 added 2011-12-22.

2000 Mathematics Subject Classification: 22D35, 22A05, 18A40.

Key words and phrases: Mackey topologies, duality, topological abelian groups.

\begin{matrix}c \\ ◯ \end{matrix} Michael Barr and Heinrich Kleisli, 2001. Permission to copy for private use granted.

commutes so that τA ⟶ A is the component at A of a natural transformation ι: τ ⟶ Id, the latter denoting the identity functor. Conversely, if we supposed that this square commutes, then evidently τf = f so that naturality of ι would give another definition of Mackey coreflections.

If SPC admits Mackey coreflections, the group A is a Mackey group if and only if τA = A. In particular, ττA = τA. If A is a Mackey group and A ⟶ B is a morphism in SPC, then we have τA = A ⟶ τB, while the continuity of τB ⟶ B gives the other half of the equality Hom(A, B) = Hom(A, τB) and thus we see that τ determines a coreflection of the inclusion SPC_{τ} ⟶ SPC.

There has been some effort into finding general conditions on a class of groups that ensure the existence of Mackey coreflections. See, for example, [Chasco et. al., 1999].

On the other hand, there has been some effort into finding a class of topological abelian groups on which the circle group is injective. It is known for locally compact groups, for example. In [Banaszczyk, 1991], especially 8.3 on page 82, it is shown that there is a large class of topological abelian groups, called nuclear groups, on which the circle group is injective. We refer the reader to the source for the definition. The class is described there as, roughly speaking, the smallest class of groups containing both locally compact abelian groups and nuclear topological vector spaces and that is closed under the operations of taking products, subgroups, and Hausdorff quotients.

What has not seemed to be noticed before (at least, as far as we can determine) is that these two questions are essentially equivalent. This is the main result of this paper, Theorem 4.1. The main tool in this construction is the so-called Chu construction, described below.

The results of this paper are in a different direction from those of Chasco et al. [1999]. In our paper, we begin with a class C of topological abelian groups and a group in that class and ask if there is a finer topology on that group that is a Mackey topology for C. They begin with a class C and a group not (necessarily) in that class and ask if there is a coarser topology on that group that is in C and is a Mackey topology for C. In particular, if the group should happen to belong to C, their conditions are conditions on which it already has a Mackey topology for C.

2. Preliminaries

The results of this section are known. See, for example, [Hewitt and Ross, 1963]. We include proofs of the simpler facts for the convenience of the reader.

We denote by T the circle group, which for our purposes it is convenient to think of as R/Z. For any n ≥ 2, we let U_{n} denote the image of the open interval (−2^{−n}, 2^{−n}) ⊆ R. These evidently form a neighborhood base at 0 for the topology. 2.1. Theorem. For any locally compact abelian group L, subgroup M ⊆ L, and character χ: M ⟶ T, there is a character ψ: L ⟶ T such that ψ|M = χ. Proof. It is shown in [Banaszczyk, 1991] that locally compact groups are nuclear (7.10) and that nuclear groups have this extension property (8.3).

We should mention that the proof that locally compact abelian groups are nuclear depends on deep theorems from [Hewitt and Ross, 1963] (see Section 23) that are used to embed any locally compact abelian group in a product C × D × E where C is compact, D is discrete and E is a finite dimensional Euclidean space.

2.2. Lemma. Let x ∈ U_{2} such that 2x ∈ U_{n} for n > 2. Then x ∈ U_{n+1}.

Proof. If 2x ∈ U_{n} then either x ∈ U_{n+1} or x − \frac{1}{2} ∈ U_{n+1}, but in the latter case, x ∉ U_{2}.

2.3. Lemma. Suppose A is a topological abelian group and χ: A ⟶ T is a homomorphism such that χ^{−1}U_{2} is open in A. Then χ is continuous.

Proof. Let V_{2} = χ^{−1}U_{2}. For each n > 2, let V_{n} be a neighborhood of 0 in A such that V_{n} + V_{n} ⊆ V_{n−1}. We claim that χ^{−1}U_{n} ⊇ V_{n}. This is true by definition for n = 2. Assuming it is true for n, let x ∈ V_{n+1}. Then 2x ∈ V_{n} and so χ(2x) = 2χ(x) ∈ U_{n} together with χ(x) ∈ χ(V_{n+1}) ⊆ χ(V_{2}) ⊆ U_{2}, gives, by Lemma 2.2, that χ(x) ∈ U_{n+1}.

2.4. Proposition. Suppose {A_{i} | i ∈ I} is a family of topological groups and A is a subgroup of ∏_{i∈I} A_{i}. Then for any character χ: A ⟶ T, there is a finite subset J ⊆ I such that χ factors through the image of A in ∏_{i∈J} A_{i}.

Proof. Suppose χ: A ⟶ T is a continuous character. Then χ^{−1}(U_{2}) is an open neighborhood of 0 in A. Thus there is a finite subset {i_{1}, …, i_{n}} ⊆ I and open neighborhoods V_{i_{1}}, …, V_{i_{n}} of A_{i_{1}}, …, A_{i_{n}}, respectively, such that

χ^{−1}(U_{2}) ⊇ A ∩ (V_{i_{1}} × ⋯ × V_{i_{n}} × ∏_{i≠i_{1},...,i_{n}} A_{i})

which implies that

χ (A ∩ (0 × ⋯ × 0 × ∏_{i≠i_{1},...,i_{n}} A_{i})) ⊆ U_{2}

and since U_{2} contains no non-zero subgroup of T,

χ (A ∩ (0 × ⋯ × 0 × ∏_{i≠i_{1},...,i_{n}} A_{i})) = 0

It follows that, algebraically at least, χ factors through the quotient

Ã = \frac{A}{A ∩ (0 × ⋯ × 0 × ∏_{i≠i_{1},...,i_{n}} A_{i})}

topologized as a subobject of A_{i_{1}} × ⋯ A_{i_{n}} by a character we will denote χ̃. To show that χ̃ is continuous, it is sufficient to observe that

χ̃^{−1}(U_{2}) ⊇ Ã ∩ (V_{i_{1}} × ⋯ × V_{i_{n}})

and invoke Lemma 2.3.

By taking A to be the whole product, we derive the following.

2.5. Corollary. The natural map ∑(A_{i} ) ⟶ (∏ A_{i}) is an isomorphism.

3. The category chu

We give a brief description here of the category chu = chu(Ab, T) on which our development depends. An object of chu is a pair (G, G′) of (discrete) abelian groups, equipped with a non-singular pairing ⟨−, −⟩: G⊗G′ ⟶ T. Non-singular means that for all x ∈ G, if x ≠ 0 there is a ϕ ∈ G′ with ⟨x, ϕ⟩ ≠ 0; similarly, for all ϕ ∈ G′, if ϕ ≠ 0 there is an x ∈ G with ⟨x, ϕ⟩ ≠ 0. It is an immediate consequence that G′ can be thought of as a subset of G and G can be thought of as a subset of G′ . Using that identification, we define a morphism f: (G, G′) ⟶ (H, H′) in chu to be a group homomorphism f: G ⟶ H such that for all ϕ ∈ H′, ϕ^{∘} f ∈ G′. Consequently, there is an induced homomorphism f′: H′ ⟶ G′ given by f′ϕ = ϕ^{∘} f. It has the property that for x ∈ G, fx is the unique element of H such that for all ϕ ∈ H′, ϕ(fx) = f′(ϕ)(x) so that f′ also determines f. Thus we can define a duality by (G, G′)^{∗} = (G′, G) with the pairing given by ⟨ϕ, x⟩ = ⟨x, ϕ⟩ for ϕ ∈ G′ and x ∈ G.

Let us denote by [(G, G′), (H, H′)] the subgroup of Hom(G, H) consisting of the morphisms described above. There is a tensor product in the category, given by

(G, G′) ⊗ (H, H′) = (G ˜⊗H, [(G, G′), (H′, H)])

This definition requires a bit of explanation. First, there is a pairing between G ⊗ H and [(G, G′), (H′, H)] given by ⟨x ⊗ y, f⟩ = ⟨y, fx⟩, for x ∈ G, y ∈ H and a morphism f: (G, G′) ⟶ (H′, H). This is non-singular in the second variable, but not in the first and we let G ˜⊗H be G ⊗ H modulo the elements of the tensor product that are annihilated by all f. Similarly, there is an internal hom given by

(G, G′) −∘(H, H′) = ([(G, G′), (H, H′)], G ˜⊗H)

The resultant category is what is called a ∗-autonomous category. This means that it has a symmetric tensor product ⊗, an internal hom −∘, natural adjunction isomorphisms

Hom(A ⊗ B, C) ≅ Hom(A, B −∘ C) ≅ Hom(B, A −∘ C)

for any objects A, B, and C; and a dualizing object ⊥ for which the natural map

A ⟶ (A −∘ ⊥) −∘ ⊥

is an isomorphism for every object A. In this case, the dualizing object is (T, Z) and one can check that (G, G′) −∘(T, Z) ≅ (G′, G). See [Barr, 1998] and references found there for further details.

For any topological abelian group A, we denote by |A| the underlying discrete group. If we suppose that each object of C has a separating family of characters, then so does each object of SPC. Then there is a functor F: SPC ⟶ chu defined on the objects by F A = (|A|, A ) with evaluation as the pairing. There is a functor R: chu ⟶ SPC defined by letting R(G, G′) be the abelian group G topologized as a subgroup of T^{G′}, topologized by the product topology. This can also be described as the weak topology for the characters in G′.

3.1. Theorem. The functor R is full and faithful and right adjoint to F.

Proof. To show that R is full and faithful, it is sufficient to show that F R ≅ Id for which it is sufficient to show that the natural map G′ ⟶ R(G, G′) is an isomorphism. But a character on R(G, G′) extends to the closure of R(G, G′) in T^{G′}, which is compact and hence, by the injectivity of T on compact groups, to T^{G′}. By Corollary 2.5, a character on T^{G′} takes the form n_{1}χ_{1} + n_{2}χ_{2} + ⋯ + n_{k}χ_{k}, where χ_{1}, χ_{2}, …, χ_{k} ∈ G′ and n_{1}, n_{2}, …, n_{k} are characters on T, that is integers. But then χ = n_{1}χ_{1} + n_{2}χ_{2} + ⋯ + n_{k}χ_{k} ∈ G′. If f: F A = (|A|, A ) ⟶ (G, G′) is given, then f: |A| ⟶ G has the property that for ϕ ∈ G′,

ϕ^{∘} f ∈ A . But this means that the composite A \underrightarrow{f} G ⟶ T^{G′} is continuous, so that f: A ⟶ R(G, G′) is continuous.

4. The main theorems

4.1. Theorem. Let C and SPC be as above and suppose that C is closed under finite products. Then the first four of the following conditions are equivalent and imply the fifth:

1. T is injective with respect to inclusions in C;

2. T is injective with respect to inclusions in SPC;

3. SPC has Mackey coreflections;

4. F has a left adjoint L whose counit LF A ⟶ A is a bijection;

5. the restriction of F to SPC_{τ} is a natural equivalence.

Proof. We will show that 1 ⟺ 2 ⟹ 3 ⟹ 4 ⟹ 2 and that 4 ⟹ 5.

1 ⟺ 2: The property is inherited by subobjects so it is sufficient to show it for products of objects of C. So suppose A ⊆ ∏_{i∈I} C_{i} with each C_{i} ∈ C. There is, by Proposition 2.4, a finite subset J ⊆ I such that χ factors by a character χ̃ through the image Ã of A in ∏_{i∈J} C_{i}. Since C is closed under finite finite products, the injectivity of T with respect to C provides the required extension of χ̃ and hence of χ. Thus 1 ⟹ 2 while the reverse inclusion is obvious.

2 ⟹ 3: Given an object A of SPC, let {A_{i} | i ∈ I} range over the set of all abelian groups whose topology is finer than that of A and for which the identity A_{i} ⟶ A induces an isomorphism A ⟶ A_{i} . Form the pullback

[figure: A; A^{I}; τA; ∏ A_{i}]

in which the bottom arrow is the diagonal and the right hand one is the identity on each factor. Since the bottom arrow is an inclusion, so is the upper one up to isomorphism. Taking duals and using the injectivity of T and Corollary 2.5, we have

[figure: A; (A^{I}) ≅ ∑_{i∈I} A; (τA); (∏ A_{i}) ≅ ∑ A_{i}; ≅]

so that A ⟶ (τA) is surjective, while it is obviously injective. Since the topology on τA is finer than that of any A_{i}, it has the finest topology that has the same character group. We must now show that τ is a functor. Suppose that f: A ⟶ B is a continuous homomorphism. Let

[figure: A; B; A′; τB]

be a pullback. Then A′ ⊆ A × τB and we have a commutative diagram (actually also a pullback)

[figure: A; A × B; A′; A × τB]

in which the top arrow (as well as the bottom) is an inclusion. Dualizing gives the commutative diagram

[figure: A; A ⊕ B; A′; A ⊕ (τB); ≅]

from which we see that A ⟶ A′ is an epimorphism and hence an isomorphism. But then A′ has a compatible topology so that τA ⟶ A′ and hence τA ⟶ τB is continuous.

3 ⟹ 4: Define L = τ ^{∘} R: chu ⟶ SPC. If f: (G, G′) ⟶ F A, we have Rf: R(G, G′) ⟶ RF A and then τR(G, G′) ⟶ τRF A. Now RF A has the same elements and same character group as A and so does τA. Since τA ⟶ A ⟶ RF A is continuous, it follows that τA has exactly the properties that characterize τRF A. Thus we have L(G, G′) ⟶ τRF A = τA ⟶ A. This gives an injection Hom((G, G′), F A) ⟶ Hom(L(G, G′), A) and the other inclusion is obvious since LF is evidently the identity.

4 ⟹ 2: Suppose that A ⊆ B in SPC. Let F A = (G, G′) and F B = (H, H′). Then G ⊆ H and we wish to show that the induced H′ ⟶ G′ is surjective. If not, let G̃′ be the image of H′ ⟶ G′. We claim that there is a pairing on (G, G̃′) such that (G, G′) ⟶ (G, G̃′) ⟶ (H, H′) are morphisms. In fact, the pairing is given by G ⊗ G̃′ ⟶ G ⊗ G′ ⟶ T. This is clearly extensional, since G̃′ is a subgroup of G′. If x ≠ 0 is in G, let ϕ ∈ H′ such that ⟨x, ϕ⟩ ≠ 0. If ϕ̃ is the image of ϕ then ⟨x, ϕ̃⟩ ≠ 0. That (G, G′) ⟶ (G, G̃′) ⟶ (H, H′) are morphisms is clear. We have the diagram

[figure: A; Ã; L(G, G′); L(G, G̃′); B; L(H, H′); R(G, G′); R(G, G̃′); R(H, H′)]

in which Ã is defined by having the lower right square be a pullback. All the vertical arrows are bijections as are the horizontal arrows in the left hand squares. Moreover, the composite A ⟶ Ã ⟶ B is the inclusion of a topological subgroup and the first arrow is a bijection, which leaves the topology of Ã both finer and coarser than that of A. Hence A = Ã, whence G′ = G̃′ and H′ ⟶ G′ is surjective.

4 ⟹ 5: If F has a left adjoint L, then it follows that T is injective in SPC and that L is constructed as above and its image is just SPC_{τ}, which is then equivalent to chu.

Suppose that A and B are topological abelian groups. Say that a homomorphism f: |A| ⟶ |B| is weakly continuous if for every continuous character ϕ: B ⟶ T, the composite ϕ^{∘} f: A ⟶ T is continuous. Clearly if every weakly continuous map out of A is actually continuous, then A is a Mackey group. The converse is also true, provided Mackey coreflections exist. We will say that C satisfies Glicksberg’s condition if every weakly continuous homomorphism between objects of C is continuous. Glicksberg proved [1962, Theorem 1.1], that the category of locally compact abelian groups does satisfy this condition.

4.2. Theorem. Suppose C satisfies Glicksberg’s condition. Then every object of C is a Mackey group in SPC.

Proof. It is sufficient to show that if C is an object of C, then for any object A of SPC, any weakly continuous f: C ⟶ A is continuous. But there is an embedding A ⟶ ∏ C_{i} with each C_{i} in C and to prove C ⟶ A continuous, it is sufficient to show that each composite C ⟶ A ⟶ C_{i} is continuous. But for any character C_{i} ⟶ T, the composite A ⟶ C_{i} ⟶ T is continuous and since C ⟶ A is weakly continuous, the composite C ⟶ A ⟶ C_{i} ⟶ T is also continuous. Thus each composite C ⟶ A ⟶ C_{i} is weakly continuous and, by Glicksberg’s condition, is continuous.

4.3. Examples. We can identify three examples, though there are certainly more. We list them in order of increasing size.

Weakly topologized groups. If we take for C the category of compact groups, the resultant SPC is the category of subcompact groups—those that have a topological embedding into a compact group. In that case, both the dual and the internal hom are topologized by the weak topology and all our results are immediately applicable. In this case the Mackey and weak topologies coincide and every object has a Mackey topology.

Locally compact groups. If for C we take the category LC of locally compact abelian groups, then SPC is the category SPLC of subobjects of products of locally compact abelian groups. Locally compact groups are Mackey groups in this case.

Nuclear groups. If we take for C the category NUC of nuclear groups, then SPC = NUC as well. Since K is injective with respect to NUC (see (8.1) in [Banaszczyk, 1991]) there are Mackey reflections. Although SPLC ⊆ NUC, it is not immediately obvious that the inclusion is proper. One point is that the class of nuclear groups is closed under Hausdorff quotients, which is not known to be the case for SPLC.

Acknowledgment

We would like to thank Jorge Galindo who brought to our attention some relevant results from the literature on topological abelian groups.

References

W. Banaszczyk (1991), Additive Subgroups of Topological Vector Spaces. Lecture Notes in Math. 1446, Springer-Verlag.

M. Barr (1998), The separated extensional Chu category. Theory and Applications of Categories 4.6, 137–147.

M. J. Chasco, E. Martín-Peinador, and V. Tarieladze (1999), On Mackey topology for groups. Studia Mathematica 132, 257–284.

I. Glicksberg (1962), Uniform boundedness for groups. Canad. Math. J. 14, 269–276.

E. Hewitt and K. A. Ross (1963), Abstract Harmonic Analysis I. Springer-Verlag.

Added 2011-12-22. It has just come to my^{1} attention that the proof that 2 ⟹ 3 in Theorem 4.1 has a gap in that we didn’t show that τ is a functor. This is easily done as follows. Suppose we have a continuous homomorphism f: B ⟶ A. Define the topological group C so that

[figure: B; A; f; C; τA; h; k; g]

is a pullback. The right hand vertical arrow and therefore the left hand vertical arrow are bijections. Suppose ϕ ∈ C . We will show that there is a ν ∈ B such that ϕ = νk, which will show that B ⟶ C is a bijection and hence that the topologyy on C lies between those of τB and B, which suffices, since then we have τB ⟶ C ⟶ τA. The definition of pullback implies that C ⟶ B × τA is a topological embedding. Injectivity of T implies that there is a (ψ, ρ) ∈ B × (τA) such that ρ = ψk + ρh. Since (τA) = A , there is a μ ∈ A such that ρ = μg. Then we have ϕ = ψk + μgh = ψk + μfk = (ψ + μf)k. Thus ν = ψ + μg is the required map.

Department of Mathematics and Statistics, McGill University 805 Sherbrooke St. W Montreal, QC H3A 2K6 Canada and Institut de Mathématiques Université de Fribourg CH-1700 Fribourg (FR) Switzerland

Email: barr@barrs.org and Heinrich.Kleisli@unifr.ch

This article may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anony-

mous ftp at ftp://ftp.tac.mta.ca/pub/tac/html/volumes/8/n4/n4.{dvi,ps}

1Unfortunately Heinrich Kleisli is no longer with us.

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi, Postscript and PDF. Details will be e-mailed to new subscribers. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh,

rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at http://www.tac.mta.ca/tac/. You may also write to tac@mta.ca to receive details by e-mail.

Editorial board.

John Baez, University of California, Riverside: baez@math.ucr.edu Michael Barr, McGill University: barr@barrs.org, Associate Managing Editor Lawrence Breen, Université Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Jean-Luc Brylinski, Pennsylvania State University: jlb@math.psu.edu Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk G. Max Kelly, University of Sydney: maxk@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Andrew Pitts, University of Cambridge: Andrew.Pitts@cl.cam.ac.uk Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca, Managing Editor Jiri Rosicky, Masaryk University: rosicky@math.muni.cz James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: walters@fis.unico.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca